If the sum of the measures of the interior angles of a polygon is twice the sum of its exterior angles how many sides does it have
Answer by jim_thompson5910(35256)

180(n-2) = 2*360

180(n-2) = 720

n - 2 = 720/180

n - 2 = 4

n = 4+2

n = 6


So there are 6 sides and this figure is a hexagon.